a line x=k intersects graph of y=log5x and the graph of  y=log5(x+4) .the distance b/w points of intersection is    0.5      , given k=a+(b)1/2..where a & b are integers..the value of    a+b  is...........

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear rahul

given distance b/w points of intersection is    0.5 

ie  log5(k+4) -log5(k)  =1/2

or log5(k+4)/k  =1/2

or (k+4)/k  =√5

  k=4/(√5 -1)

    =√5 +1

 so  a=1   and b=5

a+b=6

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin